OpenStudy (anonymous):

@jim_thompson5910 can you Help me I'm using a scale factor of 1in=ft and I measured my table which was 13ft by 8in and I 'm confused because I don't know how to do he calculation. Pleas help me @jim_thompson5910

jimthompson5910 (jim_thompson5910):

some info is missing

jimthompson5910 (jim_thompson5910):

you wrote "1in=ft " but there's no number in front of the "ft"

OpenStudy (anonymous):

Sorry it was my wall not the table

OpenStudy (anonymous):

Oh sorry 1in=1ft

jimthompson5910 (jim_thompson5910):

so every inch represents a foot

jimthompson5910 (jim_thompson5910):

if you draw a rectangle that is 1 inch long on paper then in real life, the rectangle (in this case, the wall) will be 1 foot

jimthompson5910 (jim_thompson5910):

if you have a drawing that's 8 inches by 10 inches then the real life figure will be 8 ft by 10 ft

jimthompson5910 (jim_thompson5910):

this only works because we have this 1 in = 1 ft correspondence

OpenStudy (anonymous):

So the scale factor doesn't work for 13ft by 8in

jimthompson5910 (jim_thompson5910):

is the drawing 13 ft by 8 inches? or is the real life object 13 ft by 8 inches?

jimthompson5910 (jim_thompson5910):

seems odd how you mixed ft and inches

OpenStudy (anonymous):

Yeah It is that's why I was confused it was the real life object it was my wall bu maybe it was 13ft by 8 ft is it possible

jimthompson5910 (jim_thompson5910):

13 ft by 8 ft is much better

jimthompson5910 (jim_thompson5910):

13 ft by 8 inches is way too skinny/narrow to be a wall

OpenStudy (anonymous):

Thanks so How would I do my calculation if I'm using 13ft by 8ft

jimthompson5910 (jim_thompson5910):

well if you're scaling that down to something you can draw on paper, then use the scale 1 in = 1 ft to get a 13 in by 8 in drawing

jimthompson5910 (jim_thompson5910):

Object in real life: 13 ft by 8 ft Drawing on paper: 13 in by 8 in

OpenStudy (anonymous):

Your amazing at teaching you should be a teacher you'll be the best to be honest

jimthompson5910 (jim_thompson5910):

I'm glad it's clicking

OpenStudy (anonymous):

so if 1in = 1 ft and we have 13 ft by 8 ft it'll be 13 in by 8in like you said now that 13 in by 8in could that be my dimension of scaled object.

jimthompson5910 (jim_thompson5910):

yes, your drawing will be 13 in by 8 in

jimthompson5910 (jim_thompson5910):

it will represent a real life object (the wall) that's 13 ft by 8 ft

jimthompson5910 (jim_thompson5910):

you can see that if you said something like 1 in = 1 mile, then you can draw scale models of anything really (no matter how big they are in real life)

OpenStudy (anonymous):

True and how would you show a calculation for that?is it even possible

jimthompson5910 (jim_thompson5910):

because we have this very nice ratio, we don't need to do any calculations

jimthompson5910 (jim_thompson5910):

just replace "in" with "ft" and you're done

jimthompson5910 (jim_thompson5910):

you could do (1 in)/(1 ft) = (13 in)/(x ft) 1/1 = 13/x 1*x = 1*13 x = 13 so 13 in corresponds to 13 ft but this defeats the whole convenience of the 1 in = 1 ft ratio

OpenStudy (anonymous):

so there's really no calculation to do for 13 ft by 8 ft

jimthompson5910 (jim_thompson5910):

not really

jimthompson5910 (jim_thompson5910):

so this is one instance where no math is involved and you can share this kind of info with non-math people (and not have to worry about doing calculations right or wrong)

OpenStudy (anonymous):

oh that's really cool

jimthompson5910 (jim_thompson5910):

yeah I agree